# Rethinking the Value of Agent-Generated Tests for LLM-Based Software Engineering Agents

### v1

@@ Results.
Overall, more agent-written tests do not mean more solves in this high-autonomy setting. When we push models to write more tests, test-writing status flips at scale (e.g., 64.4% for gpt-5.2), yet task outcomes remain largely unchanged: on average, 83.2% of tasks keep the same success/fail result after the intervention. At the same time, more tests can be expensive: for the low test-writing model gpt-5.2, inducing test writing increases overhead (+5.5% API calls; +19.8% output tokens) without any gain in resolution. Conversely, fewer tests can be much cheaper with only small outcome losses: for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner, suppressing test writing cuts input tokens by 49.0% and 32.9%, respectively, while success drops remain modest (2.6% and 1.8%). Taken together, varying the amount of agent-written tests strongly reshapes resource usage but has limited leverage on whether the final patch resolves the issue, suggesting that much of the test-writing effort provides low marginal utility for task resolution.
@@ 6.1. Implications
Our study indicates that the testing behaviors conducted by current LLMs, without clear guidance, do not show significant improvement in agent-based software issue resolution. It suggests two main implications for the research and practice in designing and utilizing LLM software agents.
On the one hand, our results reveal inherent limitations in how existing LLMs leverage testing for issue resolution when explicit guidance is absent. This highlights the importance of instructing agents not only on whether tests should be written, but also on how they should be designed and, more critically, how to effectively interpret and utilize the feedback generated by test execution.
On the other hand, given that current LLMs often fail to extract meaningful value from test writing, practitioners may consider adopting a more conservative approach to agent-generated tests. Such an approach could involve more sophisticated prompting strategies or guardrail mechanisms that help agents determine when test writing is truly necessary. Additionally, code agents could benefit from incorporating a cost–benefit monitoring framework that tracks the overhead associated with test-related activities, such as test writing, execution, and failure analysis, to assess their actual contribution to patch refinement.
@@ 6.2. Future work
Our findings motivate two future-work topics on agent-written testing in lightly scaffolded, high-autonomy setups.
Evaluating on-the-fly test quality in a non-stationary code state. Traditional test-quality metrics (e.g., coverage, mutation score, fault revelation) assume a fixed snapshot of the system under test and reproducible executions (Yu et al., 2023; Ryan et al., 2024; Shin et al., 2024; Bhatia et al., 2024; Yang et al., 2024b; Molinelli et al., 2025; Harman et al., 2025). In agentic development, the code state is non-stationary: tests are written and run against intermediate repository versions that may later be overwritten, and the final patch may not preserve the exact state a test originally validated. This complicates attribution (which version/function a test exercised), reproducibility, and the use of conventional quality pipelines. Future work should develop instrumentation for execution-time evaluation—e.g., capturing runnable snapshots (code, environment, commands) at test time—and define metrics that remain meaningful for transient intermediate artifacts.
Self-evolving test-generation strategies without over-constraining exploration. Our results highlight a tension between structure and autonomy: fully self-directed testing can increase cost with limited outcome gains, while a heavy, fixed workflow can narrow the search space and preclude better strategies. The goal is to produce higher-value tests by deciding when to test, what to verify (oracle design), and how to allocate budget across reasoning, editing, and validation, while preserving task-specific flexibility. A promising direction is self-evolving (Robeyns et al., 2025; Zhang et al., 2025a; Xia et al., 2025; Hu et al., 2025) testing policies: instead of a static hand-written prompt, allow the agent to revise its own testing prompt/policy from environment feedback and failure modes, adapting to task context and model capabilities (Gao et al., 2025a). Future work can formalize this as closed-loop optimization under cost and safety constraints, and compare human-specified versus self-adapted strategies under matched budgets and controlled scaffolds.
@@ 6.3. Threats to Validity
Internal validity. Agent runs can vary due to stochastic decoding and tool/environment nondeterminism, which may affect both testing behavior and resolution outcomes. Moreover, success–failure comparisons can reflect differences in task difficulty or interaction length (e.g., debugging duration), not only testing. We mitigate these concerns by treating observational results as descriptive, by using prompt-only interventions that change test creation while keeping the agent setup fixed, and by reporting task-level outcome transitions (not only aggregate resolution deltas) to show where outcomes change versus remain stable.
External validity. Our findings are based on SWE-bench under a light scaffold and a specific set of models and providers; absolute magnitudes may differ under other benchmarks, programming languages, toolchains (e.g., enforced CI), or future model versions. To support transfer, we focus on patterns that may recur in similar agent settings (e.g., large cross-model differences in testing style, observation-dominant feedback, and limited outcome sensitivity under sizable shifts in test creation), and we provide precise measurement definitions and intervention prompts to facilitate replication under alternative setups.
Data construction validity. Measurements rely on explicit operational definitions and automated extraction. Test adoption is detected through newly created test-like files, and feedback signals are extracted via deterministic AST-based rules for assertions and value-bearing prints, including common helper-style assertion APIs and a tiered taxonomy for assertion forms. These procedures can miss unconventional test artifacts, project-specific helpers, or edge-case syntax patterns. We reduce construction error by using AST parsing (rather than surface regex heuristics), conservative rules for counting value-bearing prints, and fully deterministic, reproducible extraction and categorization; nevertheless, results should be interpreted with respect to these explicit definitions.
@@ 7. Related Work
Evaluation for LLM-Generated Tests. Prior work evaluates LLM-generated testing artifacts under predefined objectives, most commonly unit tests and assertions, via systems and empirical studies on test-suite quality and model/prompt improvements (Lops et al., 2025; Yuan et al., 2024; Schäfer et al., 2023; Li et al., 2025; Yang et al., 2025), including targeted oracle generation such as assertions (Zhang et al., 2025b). Recent surveys further systematize this space, summarizing how requirements artifacts are translated into tests and the quality criteria used to judge generated tests (Yang et al., 2025). Complementing academic evaluations, industrial studies report closed-loop pipelines that combine LLM-based test generation with mutation-guided feedback to steer or refine generated tests toward stronger fault-revealing capability (Harman et al., 2025). These studies typically score outputs with fixed quality metrics (e.g., coverage, mutation-based adequacy proxies, fault revelation) on a fixed target program or code snapshot. Benchmarks similarly cast testing as a standalone objective with fixed tasks and protocols (e.g., TestEval (Wang et al., 2025), SWT-bench (Mündler et al., 2024)). In contrast, our study focuses on agent-written tests that emerge dynamically during high-autonomy, multi-step resolution of real-world GitHub issues, where the codebase and candidate patches evolve over time. We treat test writing and execution as an emergent process behavior, and characterize (i) whether/when/how intensively agents test, (ii) what signals tests encode (assertions vs. observational prints), and (iii) how these behaviors relate to resolution outcomes.
Trajectory Analysis of Software Agents. Recent work moves beyond final patch and binary success/failure by analyzing the intermediate reasoning and execution traces of LLM-based agents. Studies have examined action–observation patterns that distinguish successful from failed runs (Bouzenia and Pradel, 2025), compared trajectory length and fault-localization accuracy across agents (Majgaonkar et al., 2026), and proposed workflow taxonomies that decompose agent behavior into stages such as localization, patching, and testing-related steps (Ceka et al., 2025). Others conduct systematic failure analyses that identify root causes such as diagnostic errors and unproductive loops (Liu et al., 2025), while process-oriented studies further show that agents often hit recurrent execution errors during issue resolution, motivating lightweight checks and recovery components for robustness (Chen et al., 2026). Overall, existing trajectory analyses emphasize action sequences, outcome separation, and error categorization, but rarely examine whether and how agents autonomously decide to test. Our work addresses this gap by characterizing the emergent testing behaviors observed in execution trajectories and analyzing what feedback agent-written tests actually provide during issue resolution.
@@ 8. Conclusion
This paper revisited the common intuition that ”testing helps” for LLM-based software agents in a high-autonomy setting where writing and running tests is not specified in the prompt. Using SWE-bench Verified trajectories produced under a light agent scaffold, we separated (i) whether/when/how agents write and execute tests, (ii) what feedback those tests encode, and (iii) whether changing test-writing instructions actually shifts task outcomes and efficiency. RQ1 established a descriptive baseline of emergent testing behaviors. Across models, test writing propensity differs dramatically (from near-universal test writing to almost none), while within the same model, resolved and unresolved tasks show broadly similar test-writing rates. When tests are present, unresolved tasks tend to spread test writing across a slightly larger portion of the trajectory and execute tests more often, with higher re-execution intensity; process-level execution failures, however, vary mainly by model rather than outcome. These results highlight that in a high-autonomy setting, ”testing” is a model-specific process behavior and is not tightly coupled to eventual success. RQ2 then showed that the content of agent-written tests is largely observational. Across all models, value-revealing prints consistently outnumber assert-based checks, and overall signal volume varies widely by model. When assertions are used, their forms are broadly similar across models: most are local property checks or exact-value checks, while relational or range-style constraints are consistently rare. Taken together, agent-written tests function primarily as a runtime probing interface rather than a systematic verification mechanism. Finally, RQ3 provided controlled evidence about impact. Prompt-only interventions flip test-writing status at scale—inducing tests for the low test writing model and suppressing test writing for high test writing models—yet task outcomes are mostly unchanged (the majority of tasks preserve the same success/fail result). In contrast, efficiency effects can be substantial: for gpt-5.2—the near-zero test-writing model—encouraging test writing increases overhead without improving resolution, whereas for test-heavy models (e.g., kimi-k2-thinking and deepseek-v3.2-reasoner) discouraging test writing can sharply reduce API calls and token usage with only modest decreases in success. Overall, more agent-written tests do not mean more solves: varying test-writing effort primarily reshapes resource usage and interaction patterns, with limited leverage on whether the final patch resolves the issue.
Overall, our findings provide an empirical answer to the question: Help or Habit? In high-autonomy software issue resolution, agent-written tests often behave more like a reproduced software development lifecycle routine than a dependable source of help: models that write many tests are not consistently more successful, and prompt interventions that induce or suppress test writing rarely change whether the issue is resolved. What tests do change is the process footprint—API calls, token usage, and interaction steps—indicating that test writing frequently reflects how an agent chooses to work, not whether it can reliably validate the patch.
@@ 9. Data Availability
To support transparency and reproducibility, we publicly release a replication package containing the datasets, raw trajectories, and analysis scripts used in this study. The package is available on Figshare: https://figshare.com/s/8fa6412b86bf54fcd27e?file=61474897.

### v2

@@ Results.
Overall, varying the amount of agent-written tests strongly reshapes resource usage but has limited leverage on whether the final patch resolves the issue. More tests do not mean more solves in this high-autonomy setting, but they can impose substantial interaction overhead.
@@ 6.1. Implications
Our results suggest three practical implications. For practitioners, the goal should not be to simply make agents write more tests, but to make testing more targeted and budget-aware. One lightweight approach is to package reusable testing behaviors as focused Claude Code subagents/skills (Anthropic, 2026b, a), such as generating one minimal regression test, strengthening a weak oracle by converting prints into assertions, or running only the smallest relevant test slice. It is also useful to log test-writing, test-running, and failure-analysis costs separately so teams can see when extra test-related interaction stops adding value. For researchers, the main implication is to study testing as a process intervention, not only as a final success-rate difference. Beyond reporting outcome transitions, future studies should trace where validation budget is spent across test writing, test execution, failure inspection, and patch revision. In practice, this can be done with observability tooling such as LangSmith tracing (LangChain, 2026), making it easier to ask when testing helps, what feedback is useful, and when cheaper validation behaviors may be enough.
For benchmark users and maintainers, our study points to a narrower measurement concern. In its February 23, 2026 article “Why SWE-bench Verified no longer measures frontier coding capabilities,” OpenAI argued that SWE-bench Verified has become harder to interpret because contamination and residual benchmark-design issues weaken the meaning of final scores (OpenAI, 2026). We do not evaluate those benchmark-level claims directly, nor do we test contamination or training-data effects. Our evidence instead supports a more limited point: even when final resolution changes little, agents can still differ substantially in how often they write tests, how much validation budget they spend, and how their software-engineering behavior unfolds. This suggests that final solve rates on SWE-bench Verified are too coarse as a standalone measure of agent behavior and are better interpreted alongside process-sensitive metrics.
@@ 6.2. Future work
Two future directions follow from our results. Evaluating on-the-fly test quality in a non-stationary code state. Traditional test-quality metrics (e.g., coverage, mutation score, fault revelation) assume a fixed snapshot of the system under test and reproducible executions (Yu et al., 2023; Ryan et al., 2024; Shin et al., 2024; Bhatia et al., 2024; Yang et al., 2024b; Molinelli et al., 2025; Harman et al., 2025). In agentic development, tests are written and run against intermediate repository versions that may later be overwritten, complicating attribution and reproducibility. Future work should therefore develop execution-time instrumentation and metrics that remain meaningful for transient intermediate artifacts. Self-evolving test-generation strategies. A promising direction is self-evolving (Robeyns et al., 2025; Zhang et al., 2025a; Xia et al., 2025; Hu et al., 2025) testing policies, in which the agent revises its own testing prompt or policy from environment feedback and failure modes rather than following a static hand-written prompt (Gao et al., 2025a). Future work can formalize this as closed-loop optimization under cost and safety constraints, comparing human-specified and self-adapted strategies under matched budgets and controlled scaffolds.
@@ 6.3. Threats to Validity
Internal validity. Agent runs can vary due to stochastic decoding and tool/environment nondeterminism, which may affect both testing behavior and resolution outcomes. Moreover, success–failure comparisons can reflect differences in task difficulty or interaction length (e.g., debugging duration), not only testing. We mitigate these concerns by treating observational results as descriptive, by using prompt-only interventions that keep the agent setup fixed, and by reporting task-level outcome transitions rather than only aggregate resolution deltas.
External validity. Our findings are based on SWE-bench under a light scaffold and a specific set of models and providers; absolute magnitudes may differ under other benchmarks, programming languages, toolchains (e.g., enforced CI), or future model versions. To support transfer, we focus on patterns that may recur in similar agent settings (e.g., large cross-model differences in testing style, observation-dominant feedback, and limited outcome sensitivity under sizable shifts in test creation), and we provide precise measurement definitions and intervention prompts for replication.
Data construction validity. Measurements rely on explicit operational definitions and automated extraction. Test adoption is detected through newly created test-like files, and feedback signals are extracted via deterministic AST-based rules for assertions and value-bearing prints, including common helper-style assertion APIs and a tiered taxonomy for assertion forms. These procedures can miss unconventional test artifacts, project-specific helpers, or edge-case syntax patterns. We reduce construction error through AST parsing, conservative rules for counting value-bearing prints, and deterministic extraction and categorization; nevertheless, results should be interpreted with respect to these definitions.
@@ 7. Related Work
Evaluation for LLM-Generated Tests.
Prior work evaluates LLM-generated testing artifacts under predefined objectives, most commonly unit tests and assertions, via systems and empirical studies on test-suite quality and model/prompt improvements (Lops et al., 2025; Yuan et al., 2024; Schäfer et al., 2023; Li et al., 2025; Yang et al., 2025), including targeted oracle generation such as assertions (Zhang et al., 2025b). Recent surveys further systematize this space, summarizing how requirements artifacts are translated into tests and the quality criteria used to judge generated tests (Yang et al., 2025). Complementing academic evaluations, industrial studies report closed-loop pipelines that combine LLM-based test generation with mutation-guided feedback to steer or refine generated tests toward stronger fault-revealing capability (Harman et al., 2025). These studies typically score outputs with fixed quality metrics (e.g., coverage, mutation-based adequacy proxies, fault revelation) on a fixed target program or code snapshot. Benchmarks similarly cast testing as a standalone objective with fixed tasks and protocols (e.g., TestEval (Wang et al., 2025), SWT-bench (Mündler et al., 2024)). In contrast, our study focuses on agent-written tests that emerge dynamically during high-autonomy, multi-step resolution of real-world GitHub issues, where the codebase and candidate patches evolve over time. We treat test writing and execution as an emergent process behavior and characterize both the signals these tests encode and how those behaviors relate to resolution outcomes.
Trajectory Analysis of Software Agents.
Recent work moves beyond final patch and binary success/failure by analyzing the intermediate reasoning and execution traces of LLM-based agents. Studies have examined action–observation patterns that distinguish successful from failed runs (Bouzenia and Pradel, 2025), compared trajectory length and fault-localization accuracy across agents (Majgaonkar et al., 2026), and proposed workflow taxonomies that decompose agent behavior into stages such as localization, patching, and testing-related steps (Ceka et al., 2025). Others conduct systematic failure analyses that identify root causes such as diagnostic errors and unproductive loops (Liu et al., 2025), while process-oriented studies further show that agents often hit recurrent execution errors during issue resolution, motivating lightweight checks and recovery components for robustness (Chen et al., 2026). Overall, existing trajectory analyses emphasize action sequences, outcome separation, and error categorization, but rarely examine whether and how agents autonomously decide to test. Our work addresses this gap by characterizing emergent testing behaviors and the feedback agent-written tests provide during issue resolution.
@@ 8. Conclusion
This paper revisited the common intuition that ”testing helps” for LLM-based software agents in a high-autonomy setting where writing and running tests is not specified in the prompt. Across our three research questions, agent-written testing is better understood as a model-dependent process style than as a dependable driver of success: test-writing propensity varies sharply across models, test feedback is dominated by value-revealing prints rather than assertions, and prompt-only changes in test-writing cues usually have little effect on observed task outcomes even when they substantially change efficiency. Overall, these findings suggest that agent-written tests often behave more like a habitual software-development routine than a dependable source of validation in this setting. More agent-written tests do not mean more solves; what they more reliably change is the process footprint—API calls, token usage, and interaction patterns. Improving the value of testing for code agents may therefore require better oracles and more actionable validation signals, rather than simply inducing agents to write more tests.
@@ 9. Data Availability
We provide an anonymous Zenodo replication package containing the datasets, raw trajectories, prompt files, and analysis scripts used in this study. DOI: Phttps://doi.org/10.5281/zenodo.19251470